OpenStudy (anonymous):

if you have .0069 moles in an element how would you figure out how many atoms are in it

OpenStudy (joannablackwelder):

A mole is a number of something just like a dozen. A mole is just a lot more of that something.

OpenStudy (joannablackwelder):

.0069 moles*(6.022x10^23 atoms/1 mol) = 4.16x10^21 atoms

OpenStudy (wolfe8):

http://en.wikipedia.org/wiki/Avogadro_constant The Avogadro number is the number of atoms/molecules in 1 mol of a substance. You have 0.0069 mol so...? :)

OpenStudy (anonymous):

\[1\ mol = 6.022 \times 10^{23} \ objects\] In this case, the "objects" are your atoms. So now you can set up a conversion equation: \[(0.0069\ mol)(\frac{6.022\times10^{23}\ atoms}{1\ mol})=4.2\times10^{21}\ atoms\] Multiplying by that conversion factor is equivalent to multiplying by one, but it changes the unites of the expression!